# Self-compensating Light Calorimetry with Liquid Argon Time Projection Chamber for GeV Neutrino Physics

### v1

@@ Abstract
Liquid Argon Time Projection Chamber (LArTPC) is an exceptional dual calorimeter capable of estimating the energy of incident particles through both the ionization charge and the scintillation light. Our studies show that due to the mechanisms of charge recombination and light generation involved in the energy dissipation in liquid argon, light calorimetry in LArTPCs is inherently self-compensating: the missing energy in the hadronic component is compensated for by the extra recombination luminescence compared to the electromagnetic component. Good compensation of the electron-to-hadron response ratio ($e/h$) around unity can be achieved across a broad range of drift electric fields from 0.2 to 1.8 kV/cm. This inherent self-compensation enhances the appeal of light calorimetry in LArTPCs, complementing the well-established charge calorimetry. Using GeV neutrinos as a case study, we show that light calorimetry can achieve an energy resolution comparable to the more sophisticated charge imaging calorimetry. The synergy between light and charge calorimetry offers a novel approach to evaluating and mitigating systematic uncertainties in energy measurements with LArTPCs.
@@ I Introduction
In recent years, the technology to build massive and homogeneous liquid argon time projection chambers (LArTPCs) has advanced tremendously, owing to its many advantages in neutrino detection such as the cost-effectiveness, scalability, and high spatial resolution for particle identification. The concept was first proposed in the 1970s [15, 16] and the first large LArTPC detector, ICARUS T600 [17], operated in 2010. Many LArTPC experiments have been built since then, such as MicroBooNE [18], ICARUS [19], SBND [20], ProtoDUNE-SP [21]. and the upcoming Deep Underground Neutrino Experiment (DUNE) [22]. In the current-generation LArTPC neutrino experiments, the ionization charge signal is the main detection channel. With a uniform electric field applied in the TPC between the cathode and anode planes, the ionization electrons drift toward the anode at a constant speed. The drift time of the electron can be translated into its position along the drift direction. The transverse positions can be determined by the multiple wire planes, or printed circuit boards, at the anode where the parallel sense wires or strips are oriented at different angles. Pixelated electrodes at the anode have also been developed in recent years [23]. Together, a 3D image of the particle activities can be reconstructed with a millimeter-scale position resolution. Meanwhile, the energy deposition per unit length ($dE/dx$), also called the linear energy transfer (LET), can be measured along the particle trajectories to provide further particle identification information. Since high purity can be achieved in LAr, electrons can drift a few meters without attenuation, making it possible to build a large, fully active LArTPC calorimeter for neutrino detection.
@@
Scintillation light is used in LArTPC neutrino detectors primarily to provide the starting time ($t_{0}$) of the drift since the ionization charge signal does not provide the $t_{0}$ information by itself. Light calorimetry was not fully utilized or studied in previous LArTPC neutrino detectors primarily because of their low photon collection efficiency and large position non-uniformity. This situation may be improved in future LArTPC detectors. For instance, the proposed reference design of the DUNE Phase-II far detectors, namely, the Aluminum Profiles with Embedded X-Arapucas (APEX) design [24], aims to increase the optical surface coverage to about 60% by instrumenting the entire field cage walls with large area photon detectors. Preliminary simulations show that the average light yield of this design can reach about 180 photoelectrons per MeV, similar in performance to other organic liquid scintillator detectors used in neutrino experiments. Therefore, it is timely to revisit the LArTPC as a scintillation light calorimeter and study how well it compares and supplements the well-established ionization charge calorimetry in LArTPC.
@@
This paper is organized as follows. Section II briefly reviews the energy dissipation mechanism in LAr. In Sec. III, the deposited energy and missing energy in LArTPC is estimated for GeV neutrinos as a case study using Monte Carlo simulations. Section IV studies in detail the self-compensation mechanism observed in the LArTPC light calorimetry. Sec. V compares the energy resolution achieved from light calorimetry with charge-based calorimetry in reconstructing neutrino energy. Finally, Sec. VI discusses the applications of light calorimetry in the current and future experiments.
@@ II Energy dissipation in liquid argon
The fundamental properties of liquid argon as a particle radiation detector have been extensively studied since the 1970s. In this section, we briefly review the energy dissipation mechanisms in LAr to establish the groundwork for later sections. More details can be found in Refs. [25, 26, 27].
Energy deposition in LAr creates either excited argon atoms (Ar) or electron-ion (e-–Ar+) pairs. The ratio $\alpha$ between the number of Ar, $N_{\mathrm{ex}}$, and the number of Ar+, $N_{\mathrm{i}}$, is calculated and measured to be 0.21 [26]:
$\alpha=\frac{N_{\mathrm{ex}}}{N_{\mathrm{i}}}=0.21.$ (1)
@@
$W_{\mathrm{ph}}=\frac{W_{\mathrm{ion}}}{(1+\alpha)}=\beta\cdot W_{\mathrm{ion}}=19.5~{}\mathrm{eV},$ (2)
where $\beta=1/(1+\alpha)=0.83$. $W_{\mathrm{ph}}$ also corresponds to the energy needed to create one photon at zero electric field and will be elaborated later.
@@
$\mathrm{Ar}_{2}^{*}$$\to$ $2\mathrm{Ar}+\gamma,$
where a scintillation photon ($\gamma$) is emitted in the process. The wavelength of the emission spectrum peaks at 127 nm, in the vacuum ultra-violet (VUV) region. The collection of these scintillation photons forms the signals in light calorimetry. The number of electrons ($N_{e}$) and the number of scintillation photons ($N_{\gamma}$) created by a certain energy deposition ($E_{\mathrm{dep}}$) in LAr can be calculated as:
$N_{e}$ $=$ $N_{\mathrm{i}}=\beta\cdot\frac{E_{\mathrm{dep}}}{W_{\mathrm{ph}}},$ (4)
$N_{\gamma}$ $=$ $N_{\mathrm{ex}}=(1-\beta)\cdot\frac{E_{\mathrm{dep}}}{W_{\mathrm{ph}}}.$
For 1 MeV deposited energy, the total number of quanta created is $\sim$51000, which consists of $\sim$42000 electrons and $\sim$9000 photons. This initial distribution of quanta is illustrated in Fig. 1.
Figure 1: Illustration of the distribution of quanta between electrons ($e^{-}$) and scintillation photons ($\gamma$) before and after charge recombination in LAr. The energy deposition ($E_{\mathrm{dep}}$) is assumed to be 1 MeV and the charge recombination factor is assumed to be $R_{c}=0.7$. The number of quanta is converted into the visible energy in charge ($Q$) and light ($L$) assuming each quantum equaling 19.5 eV.
The electrons and argon ions, however, go through another process called recombination during their passage in LAr [29]:
@@
The resulting $\mathrm{Ar}^{*}$ goes through the same de-excitation process described in Eq. 3 and releases one scintillation photon. Therefore, the recombination process consumes the ionization electrons but creates more scintillation photons. The increase of scintillation light from recombination, also called recombination luminescence, was first observed in 1977 [30]. This redistribution of quanta after recombination is also illustrated in Fig. 1. It is worth noting that the total number of quanta, $N_{e}+N_{\gamma}$, is conserved before and after the recombination process. The conservation of quanta has been observed experimentally at a range of $dE/dx$ for various particles [29].
The mechanism of charge recombination in noble liquid has been studied in many literatures [31, 32]. The charge recombination factor $R_{c}$, defined as the ratio between the number of ionization electrons after recombination ($N_{e}^{r}$) and before recombination ($N_{e}$), is well described by the Birks model [31]:
$R_{c}=\frac{N_{e}^{r}}{N_{e}}=0.8/\left(1+\frac{0.0486}{\mathcal{E}\rho}\cdot\frac{dE}{dx}\right),$ (6)
@@
At zero electric field, Eq 6 indicates that all ionization electrons will recombine, and an equal amount of scintillation photons will be generated. The ideal photon yield is therefore expected to be $N_{\gamma}+N_{e}$, i.e. about 51000 photons per MeV of deposited energy at zero electric field [^1] The effects to reduce the photon yield at zero E-field, as well as other quanta reducing processes, are discussed in Appendix B. From this we can derive the light recombination factor $R_{L}$, defined as the ratio between the number of photons after recombination ($N_{\gamma}^{r}$) and the ideal photon yield, to be:
$R_{L}=\frac{N_{\gamma}^{r}}{N_{\gamma}+N_{e}}=1-\beta R_{c}.$ (7)
Figure 2 shows the charge and light recombination factors as a function of $dE/dx$ in LArTPC, assuming a drift electric field of 0.5 kV/cm, where the $dE/dx$ is normalized to that of a MIP. The recombination effect is the opposite between the charge and the light: The percentage of ionization electrons after recombination decreases for heavily ionizing particles compared with a MIP, while the percentage of scintillation photons increases. This forms the basis of the self-compensation mechanism in LArTPC light calorimetry, which will be further described in Sec. IV.
Figure 2: The recombination factor for charge ($R_{c}$) and light ($R_{L}$) as a function of $dE/dx$ in LArTPC based on Eq. 6 and 7. The external drift electric field is assumed to be 0.5 kV/cm. The $dE/dx$ is normalized to that of a minimum ionizing particle, which is about 2.1 MeV/cm.
@@ III.1 Deposited energy estimation
A LArTPC records charge and light signals as digitized waveforms. The charge sensors record induced electrical signals when ionization electrons pass by them. After dedicated signal processing and charge calibration [33], the signal waveforms can be converted to the number of ionization electrons detected, which is the same as $N_{e}^{r}$ if we neglect the detection threshold. The light sensors record the photoelectron-induced signals when optical photons strike a sensor’s photosensitive surfaces. Similarly, after signal processing and single photoelectron calibration, the signal waveforms can be converted to the number of photoelectrons ($N_{pe}$) detected, which can be further converted into the total number of optical photons ($N_{\gamma}^{r}$) given a calibrated photon collection efficiency (PCE).
There are different conventions to convert $N_{e}^{r}$ and $N_{\gamma}^{r}$ to the visible energy in charge ($Q$) and light ($L$), which only differ numerically before $Q$ and $L$ are further converted to the deposited energy $E_{\mathrm{dep}}$. To take advantage of the conservation of quanta as explained in Sec. II, we recommend that each quantum, whether of an ionization electron or a scintillation photon, represents the same amount of energy. The most convenient choice of that energy is $W_{\mathrm{ph}}$, or 19.5 eV. With this definition, $Q$ and $L$ become:
$Q=W_{\mathrm{ph}}\cdot N_{e}^{r},\quad L=W_{\mathrm{ph}}\cdot N_{\gamma}^{r}=\frac{W_{\mathrm{ph}}N_{pe}}{\mathrm{PCE}}.$ (8)
Either $Q$ or $L$ can be used to estimate the deposited energy by correcting for the recombination effect described in Sec. II:
@@
$Q+L$ $=$ $W_{\mathrm{ph}}\cdot(N_{e}^{r}+N_{\gamma}^{r})$
$=$ $W_{\mathrm{ph}}\cdot(N_{e}+N_{\gamma})$
@@ III.2 Missing energy estimation
As particles slow down, their kinetic energy eventually becomes too small to generate a detectable signal. For charge calorimetry, the detection threshold of individual energy deposition depends on the electronic noise level. In most recent LArTPCs, such as MicroBooNE [35] and ProtoDUNE-SP [36], the equivalent noise charge (ENC) has achieved below 400 $e^{-}$ owing to the implementation of cryogenic electronics. In this study, we set the charge detection threshold at 3000 $e^{-}$ for any energy deposition within 5 mm. The corresponding energy threshold is particle dependent, but it is approximately equaling to $E_{\mathrm{dep}}=100$ keV for a MIP particle with $R_{c}=0.7$. This charge threshold setting is inspired by the low energy studies in recent LArTPC experiments [37]. For light calorimetry, we assume each photon detector’s detection threshold is much below one photoelectron, so there is no loss of light signal below the threshold for GeV events.
@@
There are several other mechanisms to generate missing energy. The most common one is the charged pion and muon decays where some energy is carried away by neutrinos that escape the detector. Negatively charged pions and muons can also be captured by the Ar nuclei, resulting in some missing energy when the nuclei de-excite [39]. Neutrons will capture on Ar after thermalization and release several $\gamma$-rays that add up to about 6 MeV, which become the “negative” missing energy, i.e. extra energy if the capture $\gamma$-rays are included in the calorimetry.
@@
The number of scintillation photons after recombination ($N_{\gamma}^{r}$) is further converted to the total number of detected photoelectrons ($N_{pe}$) assuming an average absolute PCE of 0.8%. This PCE is based on a standalone Geant4 simulation study of the APEX design for DUNE Phase-II far detectors [24], where the surface optical coverage is about 60% and the photon detector (X-Arapuca) efficiency is 2%. In the current APEX design, the PCE is non-uniform across the detector volume because of the specific placements of the photon detectors. However, in this study, we assume a uniform PCE as an ideal case. Under this PCE, the light yield is about 180 photoelectrons per MeV of deposited energy for a MIP. Event-by-event fluctuation is applied to $N_{pe}$ assuming a Poisson distribution, although it does not contribute significantly to the energy resolution for GeV neutrino events at this light yield. Finally, the simulated $N_{e}^{r}$ and $N_{pe}$ are converted to the visible energy in charge ($Q$) and light ($L$) following Eq. 8.
@@
where hadrons could include protons, neutrons, charged pions, and so on, and $X$ is the nucleus after the interaction. Zero or more $\pi^{0}$s can be produced in the interaction and together with the electron from the $\nu_{e}$ interaction, they form the electromagnetic activities in the event. In the GeV neutrino energy range, the $\nu_{e}$-Ar interaction cross section consists of several components such as the quasi-elastic scattering (QES), the $\Delta$ resonance scattering (RES), the deep inelastic scattering (DIS), and so on. These interactions are modeled by the GENIE event generator to produce the final state particles and their kinematic distributions. The black histogram in Fig. 3 shows the distribution of the total available energy ($E_{\mathrm{avail}}$) of the final state particles after the $\nu_{e}$-Ar interaction. $E_{\mathrm{avail}}$ includes the kinematic energy of stable or long-lived particles such as the protons, neutrons, and electrons, and the total energy of short-lived particles including their masses such as pions and muons. Some energy is missing at the $\nu_{e}$-Ar interaction due to nuclear breakup, but the energy smearing is relatively small at the neutrino generator level.
@@
Using the simulation results such as shown in Fig. 3, we can correct for the missing energy bias by applying an additional factor to scale the mean of each distribution ($Q$, $L$, or $E_{\mathrm{dep}}$) to the initial neutrino energy $E_{\nu}$. The event-by-event fluctuation, however, cannot be corrected for and it contributes to the overall resolution of the energy reconstruction. The energy resolution is better using light calorimetry ($L$) than the simple charge calorimetry ($Q$), which will be further elaborated in Sec. IV. It is worth pointing out that the $Q+L$ method using Eq. III.1, while removing the recombination effect and recovering the $E_{\mathrm{dep}}$ distribution, does not significantly improve upon the charge-only calorimetry $Q$, because the missing energy contribution dominates the energy resolution. There are, however, other methods to further improve energy resolution in charge calorimetry by taking advantage of the imaging capability of LArTPC, which is not the main topic of this article and is described in Appendix C.

### v2

@@ Abstract
The Liquid Argon Time Projection Chamber (LArTPC) is a powerful dual calorimeter capable of estimating particle energy from both ionization charge and scintillation light. Our study shows that, due to the recombination luminescence, the LArTPC functions as a self-compensating light calorimeter: the missing energy in the hadronic component is compensated for by the increased luminescence relative to the electromagnetic component. Using 0.5–5 GeV electron neutrino charged current interactions as a case study, we show that good compensation of the electron-to-hadron response ratio ($e/h$) from 1–1.05 can be achieved across a broad range of drift electric fields (0.2–1.8 kV/cm), with better performance for neutrino energies above 2 GeV. This study highlights the potential of light calorimetry in LArTPCs for GeV neutrino energy reconstruction, complementing traditional charge calorimetry. Under ideal conditions of uniform light collection, we show that LArTPC light calorimetry can achieve an energy resolution comparable to the charge imaging calorimetry. Challenges arising from nonuniform light collection in large LArTPCs can be mitigated with a position-dependent light yield correction derived from 3D charge signal imaging.
@@ I Introduction
In recent years, the technology to build massive and homogeneous liquid argon time projection chambers (LArTPCs) has advanced tremendously, owing to its many advantages in neutrino detection such as the cost-effectiveness, scalability, and high spatial resolution for particle identification. The concept was first proposed in the 1970s [15, 16] and the first large LArTPC detector, ICARUS T600 [17], operated in 2010. Many LArTPC experiments have been built since then, such as MicroBooNE [18], ICARUS [19], SBND [20], ProtoDUNE-SP [21], and the upcoming Deep Underground Neutrino Experiment (DUNE) [22]. In the current-generation LArTPC neutrino experiments, the ionization charge signal is the main detection channel. With a uniform electric field applied in the TPC between the cathode and anode planes, the ionization electrons drift toward the anode at a constant speed. The drift time of the electron can be translated into its position along the drift direction. The transverse positions can be determined by the multiple wire planes, or printed circuit boards, at the anode where the parallel sense wires or strips are oriented at different angles. Pixelated electrodes at the anode have also been developed in recent years [23]. Together, a 3D image of the particle activities can be reconstructed with a millimeter-scale position resolution. Meanwhile, the energy deposition per unit length ($dE/dx$), also called the linear energy transfer (LET), can be measured along the particle trajectories to provide further particle identification information. Since high purity can be achieved in LAr, electrons can drift a few meters without attenuation, making it possible to build a large, fully active LArTPC calorimeter for neutrino detection.
@@
Scintillation light is used in LArTPC neutrino detectors primarily to provide the starting time ($t_{0}$) of the drift since the ionization charge signal does not provide the $t_{0}$ information by itself. Light calorimetry was not fully utilized or studied in previous LArTPC neutrino detectors primarily because of their low photon collection efficiency and large position nonuniformity. This situation may be improved in future LArTPC detectors. For instance, the proposed reference design of the DUNE Phase-II far detectors, namely, the Aluminum Profiles with Embedded X-Arapucas (APEX) design [24], aims to increase the optical surface coverage to about 60% by instrumenting the entire field cage walls with large area photon detectors. Preliminary simulations show that the average light yield of this design can reach about 180 photoelectrons per MeV, similar in performance to other organic liquid scintillator detectors used in neutrino experiments. Therefore, it is timely to revisit the LArTPC as a scintillation light calorimeter and study how well it compares to and supplements the well-established ionization charge calorimetry in LArTPC.
@@
This paper is organized as follows. Section II briefly reviews the energy dissipation mechanism in LAr. In Sec. III, the deposited energy and missing energy in LArTPC is estimated for GeV neutrinos as a case study using Monte Carlo simulations. Section IV studies in detail the self-compensation mechanism observed in the LArTPC light calorimetry. Section V compares the energy resolution achievable using light calorimetry with charge-based calorimetry in reconstructing neutrino energy. Finally, Sec. VI discusses the applications of light calorimetry in the current and future experiments.
@@ II Energy dissipation in liquid argon
The fundamental properties of liquid argon as a particle radiation detector have been extensively studied since the 1970s. In this section, we briefly review the energy dissipation mechanisms in LAr to establish the groundwork for later sections. More details can be found in Refs. [25, 26, 27, 28]. Table 1 lists the symbols used in this paper, including their values, definitions, and descriptions.
Symbol Definition Description
$\alpha$ 0.21 $N_{ex}/N_{i}$
$\beta$ 0.83 $1/(1+\alpha)$
$W_{q}$ 19.5 eV energy to create one quantum
$W_{\mathrm{ion}}$ 23.6 eV $W_{q}/\beta$
[…]
$N_{ex}$ Eq. 4 number of $\mathrm{Ar}^{*}$
$N_{q}$ $N_{i}+N_{ex}$ total number of quanta
$R_{c}$ Eq. 6 charge recombination factor
$R_{L}$ Eq. 7 light recombination factor
$N_{e}$ $N_{i}R_{c}$ electrons after recombination
$N_{ph}$ $N_{q}R_{L}$ photons after recombination
$E_{\mathrm{avail}}$ total available energy
$E_{\mathrm{dep}}$ deposited energy in LAr
$R_{\mathrm{dep}}$ $E_{\mathrm{dep}}/E_{\mathrm{avail}}$ fraction of energy not missing
$E_{\mathrm{vis}}$ visible (measured) energy
$Q$ Eq. 8 visible energy in charge
$L$ Eq. 8 visible energy in light
$E_{\mathrm{rec}}$ reconstructed energy
$R_{\mathrm{cal}}$ $E_{\mathrm{vis}}/E_{\mathrm{avail}}$ calorimeter response
$e$ EM component
$h$ hadronic component
$e/h$ $R_{\mathrm{cal}}^{e}/R_{\mathrm{cal}}^{h}$ compensation performance metric
$f_{em}$ fraction of $E_{\mathrm{vis}}$ in $e$
$\epsilon$ photon collection efficiency (PCE)
$N_{pe}$ $\epsilon N_{ph}$ number of detected photoelectrons
Table 1: List of symbols used in this paper. The symbols are organized into sections based on their relevance to quanta creation, recombination, energy reconstruction, calorimeter response, and others.
Energy deposition in LAr creates either excited argon atoms (Ar) or electron-ion (e-–Ar+) pairs. The ratio $\alpha$ between the number of Ar, $N_{ex}$, and the number of Ar+, $N_{i}$, is calculated and measured to be 0.21 [26]:
$\alpha=\frac{N_{ex}}{N_{i}}=0.21.$ (1)
@@
$W_{\mathrm{q}}=\frac{W_{\mathrm{ion}}}{(1+\alpha)}=\beta\cdot W_{\mathrm{ion}}=19.5~{}\mathrm{eV},$ (2)
where $\beta=1/(1+\alpha)=0.83$. Since $W_{q}$ also corresponds to the ideal energy needed to create one photon at zero electric field (cf. Eq. 7), it is called $W_{ph}$ in some literatures such as Ref. [26].
@@
$\mathrm{Ar}_{2}^{*}$$\to$ $2\mathrm{Ar}+ph,$
where a scintillation photon ($ph$) is emitted in the process. The wavelength of the emission spectrum peaks at 127 nm, in the vacuum ultraviolet region. The collection of these scintillation photons forms the signals in light calorimetry. The number of ionization electrons and scintillation photons created by a certain energy deposition ($E_{\mathrm{dep}}$) in LAr can be calculated as:
$N_{i}$ $=$ $\beta\cdot\frac{E_{\mathrm{dep}}}{W_{q}}$ (4)
$N_{ex}$ $=$ $(1-\beta)\cdot\frac{E_{\mathrm{dep}}}{W_{q}}$
$N_{q}$ $=$ $N_{i}+N_{ex}=\frac{E_{\mathrm{dep}}}{W_{q}},$
where $N_{q}$ is the total number of quanta, including both the ionization electrons and the scintillation photons. For 1 MeV deposited energy, $N_{q}$ is about 51000, which consists of $\sim$42000 electrons and $\sim$9000 photons. This initial distribution of quanta is illustrated in Fig. 1.
Figure 1: Illustration of the distribution of quanta between ionization electrons ($e^{-}$) and scintillation photons ($ph$) before and after charge recombination in LAr. The energy deposition ($E_{\mathrm{dep}}$) is assumed to be 1 MeV and the charge recombination factor is assumed to be $R_{c}=0.7$. The number of quanta is converted into the visible energy in charge ($Q$) and light ($L$) assuming each quantum equaling $W_{q}=19.5$ eV.
The ionization electrons and argon ions, however, go through another process called recombination during their passage in LAr [30]:
@@
The resulting $\mathrm{Ar}^{*}$ goes through the same de-excitation process described in Eq. 3 and releases one scintillation photon. Therefore, the recombination process consumes the ionization electrons but creates more scintillation photons. The increase of scintillation light from recombination, also called recombination luminescence, was first observed in 1978 by S. Kubota et al. [31], and the recombination mechanism was well articulated by A. Hitachi et al. in 1992 [32] and further elaborated by T. Doke et al. in 2002 [30]. The redistribution of quanta after recombination is illustrated in Fig. 1. It is worth noting that the total number of quanta, $N_{q}$, is conserved before and after the recombination process. The conservation of quanta in LAr has been observed experimentally at a range of LETs and electric fields for various particles [30].
The mechanism of charge recombination in noble liquid has been studied in many literatures [33, 34]. The charge recombination factor $R_{c}$, defined as the ratio between the number of ionization electrons after recombination ($N_{e}$) and before recombination ($N_{i}$), is well described by the Birks model [33]:
$R_{c}=\frac{N_{e}}{N_{i}}=0.8/\left(1+\frac{0.0486}{\mathcal{E}\rho}\cdot\frac{dE}{dx}\right),$ (6)
@@
At zero electric field, Eq. 6 indicates that all ionization electrons will recombine, and an equal amount of scintillation photons will be generated. The ideal photon yield is therefore expected to be $N_{q}$, i.e. about 51000 photons per MeV of deposited energy at zero electric field [^1] The effects to reduce the photon yield at zero E-field, as well as other quanta reducing processes, are discussed in Appendix A.. From this we can derive the light recombination factor $R_{L}$, defined as the ratio between the number of photons after recombination ($N_{ph}$) and the ideal photon yield, to be:
$R_{L}=\frac{N_{ph}}{N_{q}}=1-\beta R_{c}.$ (7)
Figure 2 shows the charge and light recombination factors as a function of $dE/dx$ in LArTPC, assuming a drift electric field of 0.5 kV/cm, where the $dE/dx$ is normalized to that of a MIP. The recombination effect is the opposite between the charge and the light: The percentage of ionization electrons after recombination decreases for heavily ionizing particles compared with a MIP, while the percentage of scintillation photons increases. The scintillation light yields in LAr for some light hadrons compared to electrons have been measured experimentally in the 1970s–1980s and summarized by T. Doke et al. [30]. These theoretical and experimental results form the basis of the self-compensation mechanism in LArTPC light calorimetry, which will be further described in Sec. IV.
Figure 2: Recombination factor for charge ($R_{c}$) and light ($R_{L}$) as a function of $dE/dx$ in LArTPC based on Eqs. 6 and 7. The external drift electric field is assumed to be 0.5 kV/cm. The $dE/dx$ is normalized to that of a minimum ionizing particle, which is about 2.1 MeV/cm.
@@ III.1 Deposited energy estimation
A LArTPC records charge and light signals as digitized waveforms. The charge sensors record induced electrical signals when ionization electrons pass by them. After dedicated signal processing and charge calibration [35], the signal waveforms can be converted to the number of ionization electrons detected, which is the same as $N_{e}$ if we neglect the detection threshold. The light sensors record the photoelectron-induced signals when optical photons strike a sensor’s photosensitive surfaces. Similarly, after signal processing and single photoelectron calibration, the signal waveforms can be converted to the number of photoelectrons ($N_{pe}$) detected, which can be further converted into the total number of optical photons ($N_{ph}$) given a calibrated photon collection efficiency (PCE) $\epsilon$: $N_{ph}=N_{pe}/\epsilon$.
There are different conventions to convert $N_{e}$ and $N_{ph}$ to the visible energy in charge ($Q$) and light ($L$), which only differ numerically before $Q$ and $L$ are further converted to the deposited energy $E_{\mathrm{dep}}$. To take advantage of the conservation of quanta as explained in Sec. II, we recommend that each quantum, whether of an ionization electron or a scintillation photon, represents the same amount of energy. The most convenient choice of that energy is $W_{q}$, or 19.5 eV. With this definition, $Q$ and $L$ become:
$Q=W_{q}\cdot N_{e},\quad L=W_{q}\cdot N_{ph}.$ (8)
Both $Q$ and $L$ can be used to estimate the deposited energy by correcting for the recombination effect described in Sec. II:
@@
$Q+L$ $=$ $W_{q}\cdot(N_{e}+N_{ph})$
$=$ $W_{q}\cdot(N_{i}+N_{ex})$
@@ III.2 Missing energy estimation
As particles slow down, their kinetic energy eventually becomes too small to generate a detectable signal. For charge calorimetry, the detection threshold of individual energy deposition depends on the electronic noise level. In most recent LArTPCs, such as MicroBooNE [37] and ProtoDUNE-SP [38], the equivalent noise charge has achieved below 400 $e^{-}$ owing to the implementation of cryogenic electronics. In this study, we set the charge detection threshold at 3000 $e^{-}$ for any energy deposition within 5 mm. The corresponding energy threshold is particle dependent, but it is approximately equal to $E_{\mathrm{dep}}=100$ keV for a MIP particle with $R_{c}=0.7$. This charge threshold setting is inspired by the low energy studies in recent LArTPC experiments [39]. For light calorimetry, we assume each photon detector’s detection threshold is much below one photoelectron, so there is no loss of light signal below the threshold for GeV events.
@@
There are several other mechanisms to generate missing energy. The most common one is the charged pion and muon decays where some energy is carried away by neutrinos that escape the detector. Negatively charged pions and muons can also be captured by the Ar nuclei, resulting in some missing energy when the nuclei de-excite [41]. Neutrons will capture on Ar after thermalization and release several $\gamma$ rays that add up to about 6 MeV, which become the “negative” missing energy, i.e. extra energy if the captured $\gamma$ rays are included in the calorimetry.
@@
The number of scintillation photons after recombination ($N_{ph}$) is further converted to the total number of detected photoelectrons ($N_{pe}$) assuming an average absolute PCE of 0.8%. This PCE is based on a standalone GEANT4 simulation study of the APEX design for DUNE Phase-II far detectors [24], where the surface optical coverage is $\sim$60and the photon detector (X-Arapuca) efficiency is $\sim$2%. In the current APEX design, the PCE is nonuniform across the detector volume because of the specific placements of the photon detectors. However, in this study, we assume a uniform PCE as an ideal case. Under this PCE, the light yield is about 180 photoelectrons per MeV of deposited energy for a MIP. Methods to mitigate nonuniform light collection are discussed in Sec. V.2. Event-by-event fluctuation is applied to $N_{pe}$ assuming a Poisson distribution, although it does not contribute significantly to the energy resolution for GeV neutrino events at this light yield. Finally, the simulated $N_{e}$ and $N_{pe}$ are converted to the visible energy in charge ($Q$) and light ($L$) following Eq. 8.
@@
where hadrons could include protons, neutrons, charged pions, and so on, and $X$ is the nucleus after the interaction. Zero or more $\pi^{0}$s can be produced in the interaction and together with the electron from the $\nu_{e}$ interaction, they form the electromagnetic activities in the event. In the GeV neutrino energy range, the $\nu_{e}$-Ar interaction cross section consists of several components such as the quasielastic scattering, the $\Delta$ resonance scattering, the deep inelastic scattering, and so on. These interactions are modeled by the GENIE event generator to produce the final state particles and their kinematic distributions. The black histogram in Fig. 3 shows the distribution of the total available energy ($E_{\mathrm{avail}}$) of the final state particles after the $\nu_{e}$-Ar interaction. $E_{\mathrm{avail}}$ includes the kinematic energy of stable or long-lived particles such as the protons, neutrons, and electrons, and the total energy of short-lived particles including their masses such as pions and muons. Some energy is missing at the $\nu_{e}$-Ar interaction due to nuclear breakup, but the energy smearing is relatively small at the neutrino generator level.
@@
Figure 4: (Top) Energy distribution in a $\nu_{e}$-Ar CC interaction. The pie chart shows how the total available energy ($E_{\mathrm{avail}}$) in a $\nu_{e}$-Ar CC interaction is divided, based on an average of $10^{4}$ simulated 3 GeV $\nu_{e}$ CC events (cf. Eq. 11). The $E_{\mathrm{avail}}$ is allocated across various categories of primary particles’ deposited energy (protons, neutrons, charged pions, EM components, and others) as well as the total missing energy. For each primary particle, its deposited energy includes contributions from both the particle itself and its secondary particles as they interact with LAr. (Bottom) For each category of primary particles, their available energy is further divided into deposited energy and missing energy. The respective energy ratios are displayed for each particle category.
Figure 4 further illustrates the missing energy contribution in $\nu_{e}$-Ar CC interactions. The pie chart in the top panel shows how the total available energy ($E_{\mathrm{avail}}$) in a $\nu_{e}$-Ar CC interaction is divided, based on an average of $10^{4}$ simulated 3 GeV $\nu_{e}$ CC events (cf. Eq. 11). The $E_{\mathrm{avail}}$ is allocated across various categories of primary particles and the total missing energy. For each primary particle, its deposited energy includes contributions from both the particle itself and its secondary particles as they interact with LAr. On average, 64.5% of $E_{\mathrm{avail}}$ is distributed in the EM component, consisting of electrons and $\pi^{0}$s. The hadronic component is divided into protons, neutrons, charged pions, and other particles such as kaons and lambda baryons, where their deposited energy over $E_{\mathrm{avail}}$ are 10.4%, 6.0%, 9.8%, and 1.3%, respectively. The total missing energy that is unmeasured is 8.0% on average. In the bottom panel, the available energy of each primary particle after a $\nu_{e}$-Ar CC interaction is divided into its deposited energy and missing energy. The average percentage of missing energy is 11.8%, 35.8%, 24.7%, and 0.04% for protons, neutrons, charged pions, and the EM component, respectively, indicating that missing energy mostly comes from the hadronic component.
@@
Using the simulation results such as shown in Fig. 3, we can correct for the missing energy bias by applying an additional factor to scale the mean of each distribution ($Q$, $L$, or $E_{\mathrm{dep}}$) to the initial neutrino energy $E_{\nu}$. The event-by-event fluctuation, however, cannot be corrected for and it contributes to the overall resolution of the energy reconstruction. The energy resolution is better using light calorimetry ($L$) than the simple charge calorimetry ($Q$), which will be further elaborated in Sec. IV. It is worth pointing out that the $Q+L$ method using Eq. III.1, while removing the recombination effect and recovering the $E_{\mathrm{dep}}$ distribution, does not significantly improve upon the charge-only calorimetry $Q$, because the missing energy contribution dominates the energy resolution. There are, however, other methods to further improve energy resolution in charge calorimetry by taking advantage of the imaging capability of LArTPC, which is not the main topic of this article and is described in Appendix B.